OpenStudy (anonymous):

find the inverse of H(x)= (X+3)^2 with restrictions of ( negative infinity, -3]

OpenStudy (abmon98):

H(x)=(x+3)^2 y=(x+3)^2 x=(y+3)^2 (x)^1/2=y+3 y=(x)^1/2-3

OpenStudy (abmon98):

f^-1(x)=(x)^1/2-3

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[x=(y+3)^2\] solve for \(y\)

OpenStudy (anonymous):

start with \[\pm\sqrt x=y+3\] then \[y=\pm\sqrt{x}-3\] but since you are on \((-\infty, 3\) it is \[f^{-1}(x)=-\sqrt{x}-3\]

OpenStudy (anonymous):

@satellite73 thank you so much!!